### v1

# ReviewAgents: Bridging the Gap Between Human and AI-Generated Paper Reviews

Abstract
Academic paper review is a critical yet time-consuming task within the research community. With the increasing volume of academic publications, automating the review process has become a significant challenge. The primary issue lies in generating comprehensive, accurate, and reasoning-consistent review comments that align with human reviewers’ judgments. In this paper, we address this challenge by proposing ReviewAgents, a framework that leverages large language models (LLMs) to generate academic paper reviews. We first introduce a novel dataset, Review-CoT, consisting of 142k review comments, designed for training LLM agents. This dataset emulates the structured reasoning process of human reviewers—summarizing the paper, referencing relevant works, identifying strengths and weaknesses, and generating a review conclusion. Building upon this, we train LLM reviewer agents capable of structured reasoning using a relevant-paper-aware training method. Furthermore, we construct ReviewAgents, a multi-role, multi-LLM agent review framework, to enhance the review comment generation process. Additionally, we propose ReviewBench, a benchmark for evaluating the review comments generated by LLMs. Our experimental results on ReviewBench demonstrate that while existing LLMs exhibit a certain degree of potential for automating the review process, there remains a gap when compared to human-generated reviews. Moreover, our ReviewAgents framework further narrows this gap, outperforming advanced LLMs in generating review comments.
[…]
ReviewAgents: Bridging the Gap Between Human and AI-Generated Paper Reviews
Xian Gao1, Jiacheng Ruan1, Jingsheng Gao1, Ting Liu1, Yuzhuo Fu1 1Shanghai Jiao Tong University
1 Introduction
Peer review of academic papers is a crucial component of the scholarly publishing system, as it ensures the quality of scientific research and facilitates the improvement of academic writing. However, with the rapid increase in the number of academic paper submissions, the traditional peer review process is confronted with challenges such as inefficiency and a limited number of reviewers, resulting in restricted feedback for authors and hindering the timely acquisition of review comments.
The reasoning capabilities of LLMs for complex tasks render them potentially valuable for providing feedback on academic papers Liu and Shah (2023); Zhao et al. (2024); Zhuang et al. (2025). This, to some extent, alleviates the pressure on reviewers and allows authors to independently use LLMs to review and revise their papers prior to formal submission. However, existing research and applications primarily focus on prompting LLMs to directly generate comments on the submitted papers Zhou et al. (2024); D’Arcy et al. (2024), which greatly simplifies the review process. This approach does not align with the cognitive process of human reviewers in real-world peer review, nor with the multi-step, multi-role nature of the peer review process. As a result, comments generated directly by LLMs are often difficult to align with the review comments provided by human reviewers. In fact, when human reviewers write their reviews, they are typically required to first summarize the main content of the paper, followed by enumerating its strengths and weaknesses, and ultimately reaching a review conclusion. After multiple reviewers have written their comments, the area chair (AC) synthesizes the feedback from all reviewers and provides a final summary as a meta-review. This process ensures that each reviewer fully understands the paper’s content, guaranteeing the accuracy, objectivity, and fairness of the review outcomes while minimizing the influence of individual reviewer biases on the final review comments.
To bridge the gap between LLM-based reviews and human review behavior, we have developed a multi-agent review framework, named ReviewAgents, which emulates the multi-step reasoning process of human reviewers. To train our agents, we collected a substantial amount of review comments from publicly available review platforms and structurally transcribed them, constructing the Review-CoT dataset, which guides the training of both reviewer agents and area chair agents. Specifically, Review-CoT contains 37,403 papers and 142,324 corresponding review comments and meta-reviews, collected from open peer review platforms. We utilized state-of-the-art large language models to transcribe the review comments into a stepwise reasoning format, aligning them with the cognitive process of human reviewers. Considering the timeliness of novelty assessments in review comments, we incorporated references to the most relevant papers, published up to the point of submission, into the dataset, and employed a relevant-paper-aware training method. This approach mirrors the process by which human reviewers evaluate novelty by retrieving relevant papers. The dataset is used to train both the reviewer and area chair agents. Upon completion of training, we employ ReviewAgents, a multi-agent framework, to simulate the review process by human reviewers and area chairs, utilizing multi-agent, multi-step review procedures to generate the final review comments. Unlike previous approaches, our ReviewAgents framework leverages the knowledge of relevant papers and fully emulating the paper evaluation process conducted by human experts.
To comprehensively assess the quality of generated review comments and their alignment with human reviewers’ feedback, we propose ReviewBench, a benchmark specifically designed to evaluate the quality of automatically generated review comments. ReviewBench includes the latest papers and review comments from open platforms, ensuring that its data is not included in the pretraining datasets of most existing LLMs. We evaluate the capabilities of large language models in generating paper review comments across four dimensions.
The main contributions of our work are summarized as follows:
  1. 1.
    We introduce the Review-CoT dataset, which, to the best of our knowledge, is the first and largest dataset to simulate the human review thought process and includes relevant papers.
  2. 2.
    We propose ReviewBench, a benchmark for the quantitative evaluation of review comments generated by LLMs.
  3. 3.
    By employing a relevant-paper-aware training method on the Review-CoT dataset, we construct ReviewAgents, a multi-step, multi-role framework for paper review that simulates the human review process. Experimental results demonstrate that our approach generates review comments closely aligned with those of human reviewers.
2 Related Work
2.1 Large Language Models for Paper Review
Large language models (LLMs) have demonstrated significant potential in reviewing and comprehending complex texts Liu and Shah (2023); Zhao et al. (2024); Zhuang et al. (2025); Ruan et al. (2025); Zhu et al. (2024); Ruan et al. (2024b, a). Initial studies, which compared reviews generated by human reviewers and LLMs for academic papers, indicated that LLMs can produce comments with substantial overlap with those of human reviewers, contributing effectively to the peer review process Robertson (2023); Liang et al. (2023). However, further research has shown that, despite the advanced capabilities of LLMs such as GPT-3.5 and GPT-4 in scoring input paper texts, the generated reviews still fall short of fully meeting human expectations Zhou et al. (2024). Researchers have fine-tuned LLMs to accommodate review requirements by constructing datasets from publicly available review comments Kang et al. (2018); Yuan et al. (2021); Shen et al. (2022); Dycke et al. (2023); Gao et al. (2024). Additionally, some studies have employed multi-turn dialogues Tan et al. (2024) or prompted multiple LLMs to construct review systems D’Arcy et al. (2024), providing comprehensive feedback on the entire text. However, unlike current studies that focus solely on generating review comments, our work aims to simulate the entire thought process and review procedure across the full peer review process. We emphasize the individual deep thinking and multi-step, multi-role nature of human peer review in the real world. This approach alleviates biases inherent in reviews generated by a single LLM, thereby aligning more closely with human-generated review feedback.
2.2 Chain-of-thought Reasoning in Large Language Models
Through chain-of-thought prompting Wei et al. (2022), LLMs are guided with a step-by-step reasoning process, enabling them to break down complex problems into a series of incremental reasoning steps Chu et al. (2024). This decomposition enhances LLMs’ performance in reasoning tasks, advancing structured prompting techniques Qiao et al. (2024); Cesista (2025). Recent studies highlight that the systematization and structuring of reasoning significantly affect LLMs’ performance, particularly their ability to perform multi-stage reasoning and identify key tasks at each stage. This can be achieved through independent language reasoning systems Zhong et al. (2024) or supervised fine-tuning Xu et al. (2025). In this paper, we introduce structured reasoning to automated peer review, aligning LLM behavior with human review practices by dividing the review process into a three-stage structured procedure, improving review generation outcomes.
2.3 Instruction Tuning Dataset
Instruction fine-tuning is a specialized training process for LLMs designed to improve their ability to follow specific instructions and perform tasks with higher precision and reliability. This process relies on a dataset containing task-specific examples paired with clear instructions. For academic paper review tasks, various datasets of review comments have been developed, such as those for the NLP community Kang et al. (2018); Dycke et al. (2023) and datasets spanning multiple research domains Gao et al. (2024). Some datasets also include more detailed content, such as meta-review comments Shen et al. (2022) or multi-turn dialogues to demonstrate the rebuttal process Tan et al. (2024). However, these datasets do not capture the reviewer’s thought process. We introduce Review-CoT and ReviewBench, which together establish the first pipeline for LLM-driven automated paper review, encompassing both upstream training and downstream evaluation. To our knowledge, this is the first dataset to structurally represent the review steps.
3 Dataset and Benchmark
3.1 Structured Thinking in Review
Given that paper review requires models to thoroughly read the manuscript, integrate external knowledge, and conduct an in-depth analysis of the paper’s strengths and weaknesses before assigning scores, a detailed stepwise reasoning process is essential. Inspired by the cognitive process of human reviewers, we decompose the review process of the intelligent review agent into three structured stages: Summarization, Analysis, and Conclusion, thereby aligning the automated review procedure with human-like reasoning.
  • Summarization Stage: The agent is provided with the full text of the paper and is tasked with concisely summarizing its core content and methodologies.
  • Analysis Stage: The agent evaluates the strengths and weaknesses of the paper by considering its content in conjunction with relevant works in the same domain. This analysis serves as the foundation for subsequent assessments.
  • Conclusion Stage: Based on the insights gained from the previous two stages, the agent formulates a final review decision for the paper.
To facilitate this structured stepwise reasoning, we employ dedicated tags (e.g., <SUMMARY>…</SUMMARY>) to explicitly mark the beginning and end of each stage. These tags enable the model to maintain clarity throughout the entire reasoning process.
3.2 Dataset Preparation
We have defined a standardized data format for Review-CoT. Each data entry contains the following key information: a unique identifier, paper title, abstract, full-text content, submission conference, reviewer comments, Program Chair’s remarks, and the decision regarding paper acceptance. Additionally, the data records the titles and abstracts of the most relevant papers in the field up to the point of submission. With this information, the dataset samples we construct can be aligned as closely as possible with the information typically accessed by human reviewers during the review process, enabling the trained models to generate conclusions that are more consistent with those of human reviewers.
Figure 1: The process of transcribing the collected review comments using advanced LLMs according to a pre-defined structured thinking process
3.2.1 Metadata Collection
We crawled ICLR papers from 2017 to 2024 and NeurIPS papers from 2016 to 2024 from OpenReview [^1] https://openreview.net/ and NeurIPS Proceedings [^2] https://papers.nips.cc/. For each paper, we adhere to the standardized data format we have defined, retaining as much metadata information as possible. Each paper includes the full text as well as the corresponding review comments. For the paper content, we use Scipdf_Parser [^3] https://github.com/titipata/scipdf_parser to parse the PDFs into structured JSON files.
3.2.2 Transcription
An examination of review comments within the collected metadata revealed that recent comments already follow a predefined structured review process, including a paper summary, analysis of strengths and weaknesses, and a final conclusion. While earlier comments contained these elements, they lacked explicit structural divisions. To address this, we employed advanced LLMs to transcribe the review comments according to the structured reasoning process (Figure 1). This transcription reorganizes the paragraphs while preserving the original content and semantics, adding dedicated tags (e.g., <SUMMARY>…</SUMMARY>). This process transforms the review comments into a unified format suitable for model training.
3.2.3 Relevant Paper Retrieval
Human reviewers typically reference relevant literature to understand the field’s current state to assess a paper’s novelty, ensuring the timeliness of their review comments. The review comments in Review-CoT dataset, which evaluate content and novelty, reflect the reviewer’s judgments during the review process. These judgments must be compared with the latest research to maintain timeliness, as outdated reviews cannot be generalized across time without considering recent findings. To enhance the reviewer’s thought process and ensure timely comments, we use the Semantics Scholar API [^4] https://www.semanticscholar.org/product/api to retrieve papers in the same research domain as the reviewed paper, up to the submission date. We extract titles and abstracts from these papers and incorporate them into the dataset as references to assess the paper’s novelty. Both the paper itself and the titles and abstracts of the retrieved relevant papers are encoded as vectors, and their similarity is computed. We retain the two papers with the highest similarity as the most relevant references for the reviewer during the review process.
Figure 2: An overview of the ReviewAgents framework
3.2.4 Dataset Details
We have collected a total of 142,324 review comments for 37,403 papers from publicly available review platforms. Each review comment was transcribed into a structured format with appropriate delimiters added. For each paper, we included the titles and abstracts of the two most relevant papers, published up to the submission date, as references for the review process.
For the training dataset of the Review Agent, the model’s input consists of the original paper and the two relevant papers, with the output being the formatted review comments. For the Area Chair Agent’s training dataset, the model’s input is the output from the Review Agent, and the output is the paper’s meta-review.
Table 1 presents a comparison between our dataset and existing review comment datasets, showing that our dataset outperforms the existing ones in both data volume and diversity.
# Papers # Reviews
Meta
Review
Thinking
Process
Relevant
Papers
PeerRead
Kang et al. (2018)
3,006 10,770
ASAP-Review
Yuan et al. (2021)
8,877 28,119
MReD
Shen et al. (2022)
7,894 30,764
NLPeer
Dycke et al. (2023)
5,672 11,515
Review2
Gao et al. (2024)
27,805 99,729
Ours 37,403 142,324
Table 1: Comparison between Review-CoT and existing review comment datasets
3.3 ReviewBench
Given that the evaluation of LLM-generated review comments is still limited to model scoring and human ratings, with a lack of standardized quantitative assessment criteria, we propose ReviewBench, a benchmark for evaluating LLM-generated review comments across multiple dimensions. Table 2 presents the evaluation dimensions and corresponding metrics included in ReviewBench. Our benchmark encompasses assessments of LLM-generated review comments in terms of language quality, semantic consistency, and sentiment consistency, as well as a Review Arena task to measure the model’s ability to generate review comments. Further details on the evaluation metrics can be found in Appendix.
3.3.1 Language Diversity
[…]
3.3.2 Semantic Consistency
Semantic consistency is employed to assess the degree of semantic alignment between the review comments generated by the model and those written by human reviewers, using the ROUGE-1, ROUGE-L, and SPICE metrics. ROUGE Lin (2004) is a widely used metric for evaluating the overlap between generated text and reference text. ROUGE-1 measures the overlap of individual words between the generated and reference texts, while ROUGE-L evaluates the longest common subsequence (LCS) between them. The SPICE metric Anderson et al. (2016), originally designed to evaluate the consistency of image captions, measures semantic consistency by comparing part-of-speech and semantic structures, assessing whether the generated text aligns semantically with the reference text.
3.3.3 Sentiment Consistency
Sentiment consistency is employed to measure the overall alignment of sentiment and attitudinal orientation between the review comments generated by the model and those written by human reviewers, specifically assessing whether both are predominantly positive or negative Wang and Wan (2018). The sentiment tendencies of both model-generated and human review comments are extracted using sentiment analysis models based on BERT Devlin et al. (2019) and VADER Hutto and Gilbert (2014), respectively. The distance between these tendencies is then evaluated as the sentiment consistency scores, BERTScore and VADER Score.
3.3.4 Review Arena
The open nature of peer review presents significant challenges for automated evaluation. Previous studies have shown that direct scoring methods using LLMs do not align well with human preferences Liang et al. (2023). In contrast, LLMs perform better in ranking tasks Li et al. (2024). To obtain reliable evaluations, we propose Review Arena, which employs a tournament-style pairwise evaluation system to compare pairs of review comments. To mitigate the influence of hallucination, we use human review comments as a reference and require LLM judges to rank a pair of model-generated review comments based on their similarity in content and form to the human review comments. Each pair of review comments undergoes two evaluations through reverse sorting to minimize positional bias.
Dimensions Metrics
Language Diversity $\text{Distinct}_{4}$
Inverse Self-BLEU@4
Semantic Consistency ROUGE-1
ROUGE-L
SPICE
Sentiment Consistency BERTScore
VADER Score
Review Arena Win Rate
Table 2: Evaluation dimensions and corresponding metrics in ReviewBench
3.3.5 Statistical Information
We selected 100 papers from ICLR 2024 and NeurIPS 2024 as the test set for our benchmark. ICLR 2024 and NeurIPS 2024 were published in February and September of 2024, respectively, both of which postdate the knowledge cutoff times of most LLMs’ pretraining datasets [^5] The knowledge cutoff time for the GPT-4o and Llama-3.1 is October 2023 and December 2023, respectively., ensuring that the test data was not used during the pretraining phase. These papers were removed from the Review-CoT training dataset. To better simulate the real review process, the selected benchmark follows an acceptance-to-rejection ratio of 3:7, which closely approximates the acceptance rate of these two conferences.
Model Language Diversity $\uparrow$ Semantic Consistency $\uparrow$
Sentiment
Consistency
$\uparrow$
Overall $\uparrow$
$\text{Distinct}_{4}$
Inverse
Self-BLEU@4
ROUGE-1 ROUGE-L SPICE
BERT
Score
VADER
Score
Human 98.88 89.84 100.00 100.00 100.00 100.00 100.00 98.39
API-based LLMs
[…]
Deepseek-R1 99.37 77.64 37.22 14.40 13.57 48.13 83.09 53.35
[…]
Deepseek-V3 95.51 69.36 38.06 16.57 14.58 30.20 85.70 50.00
[…]
Open-source LLMs
Llama-3.1-70B-Instruct 87.58 60.94 40.53 17.83 15.24 35.56 84.02 48.81
Qwen-2.5-72B-Chat 94.72 64.93 29.97 15.56 15.18 26.42 84.66 47.35
Llama-3.1-8B-Instruct 82.98 55.70 30.39 13.90 12.95 26.23 81.00 43.31
[…]
Table 3: Comparison of the diversity and consistency metric scores
4 Methodology
4.1 Framework Overview
Figure 2 illustrates an overview of the ReviewAgents framework we propose. During the training phase, we employ a training strategy that incorporates awareness of relevant papers and a structured reasoning approach to train the reviewer agent, while using meta-review to train the area chair agent, enabling the model to simulate the human peer review process. In the inference phase, the reviewer, trained with awareness of relevant papers, generates review comments by considering relevant papers retrieved from the literature database. The area chair then synthesizes multiple review comments to form the final review conclusion.
4.2 Relevant-paper-aware Training
The timeliness of review comments is reflected in the relevant papers that have been published up until the time of the review. In the Review-CoT training dataset, relevant papers are provided as references for the review process. During training, the model learns how to integrate these relevant papers into the generation of review comments, thereby incorporating timeliness information and closely replicating the reasoning process employed by human reviewers when crafting their feedback. To ensure that the model accesses the relevant paper information without exceeding the context window, we input the titles and abstracts of the relevant papers as prompts during training.
4.3 Multi-step Multi-role Review Generation
After training LLM agents for different roles, we construct a review framework that emulates the human review process. The paper is initially assigned to multiple distinct reviewers, each of whom utilizes a retrieval API to search for relevant papers in the literature database as references for their review. Each reviewer then follows a structured reasoning process in sequence—SUMMARY, ANALYZE, and CONCLUDE—to generate their review comments. Subsequently, the review comments are submitted to the area chair agent for aggregation and synthesis, eliminating any biases present in the individual reviewer’s feedback. The final review comments, provided by the area chair, serve as the ultimate output of the review process.
Figure 3: Win rates of ReviewAgents in Review Arena
5 Experiments
5.1 Experimental Setup
We evaluated our method on ReviewBench, selecting advanced open-source models and closed-source models for comparison, evaluating the meta-reviews generated by different models. The evaluation metrics, Distinct and Inverse Self-BLEU, are computed using 4-grams, denoted as $\text{Distinct}_{4}$ and Inverse Self-BLEU@4, respectively. For ROUGE-1 and ROUGE-L, their F1-scores are computed. For the diversity and consistency metrics, including Language Diversity, Semantic Consistency, and Sentiment Consistency, we compute their overall scores.
Our framework utilizes the Llama-3.1-8B-Instruct model trained on NVIDIA A100 GPUs. When generating review comments, we set the number of reviewers, $N$, to 3. The judge for the Review Arena is GPT-4o. Detailed implementation specifics and prompt templates are provided in Appendix.
5.2 Main Results
5.2.1 Diversity and Consistency
Table 3 presents a comparison of the diversity and consistency metric scores of our ReviewAgents with those of other models on ReviewBench. It is evident that, despite being constrained by the model’s parameter size, the proposed ReviewAgents method exhibits slight limitations in language diversity. However, the review comments generated by ReviewAgents outperform larger parameter open-source models in both semantic consistency and viewpoint alignment, achieving results comparable to or exceeding those of advanced closed-source models. In terms of overall metrics, our approach achieves state-of-the-art results, validating the comprehensiveness and effectiveness of our method.
$N$ of Reviewers Language Diversity $\uparrow$ Semantic Consistency $\uparrow$
Sentiment
Consistency
$\uparrow$
Overall
Distinct4
Inverse
Self-BLEU@4
ROUGE-1 ROUGE-L SPICE
BERT
Score
VADER
Score
1 97.78 82.29 43.20 19.41 15.66 31.67 84.83 53.55
[…]
5 96.16 75.95 42.23 19.04 15.55 42.98 85.85 53.97
6 95.81 75.24 42.25 19.08 15.67 43.39 85.91 53.91
Table 4: Ablation study on the number of reviewer agents
5.2.2 Review Arena Results
Figure 3 illustrates the win rates of ReviewAgents compared to other models in the Review Arena. ReviewAgents achieved a higher win rate than closed-source LLMs and significantly outperformed open-source LLMs, further validating the effectiveness of our approach in generating review comments that align closely with human feedback.
5.3 Ablation Study on Number of Reviewer Agents
The number of agents in multi-agent systems directly affects the final outcome Zhang et al. (2024). To determine the optimal number of reviewers within the ReviewAgents framework, we examined the performance of review comments generated with different reviewer counts. Table 4 shows the impact of varying values of $N$ on the generated review comments. The results indicate that when $N=3$, the generated review comments align most closely with human reviewers’ feedback. Further experiments reveal that fewer reviewers yield more diverse comments, while an excessive number of reviewers does not improve alignment with human opinions. This may be due to the fact that, in the training dataset, the average number of review comments per paper is between 3 and 4. As a result, too many review comments during inference introduce opinion discrepancies and extended context, negatively affecting the area chair agent’s judgment.
5.4 Discussion and Case Study
As shown in Table 3, the improvements of ReviewAgents over other methods primarily stem from enhanced semantic and sentiment similarity with human review comments. This improvement is primarily due to the structured generation process of review comments, which mirrors the stages of human peer review. As a result, each LLM stage generates review comments through a reasoning process closely aligned with human reviewers. Additionally, the final meta-review is derived from aggregating comments from multiple reviewer agents, reducing individual reviewer biases.
Figure 4 presents a case study. Compared to comments from advanced LLMs, ReviewAgents’ comments show greater overlap with human reviewers’ comments. For example, both address issues such as insufficient comparison and citation, highlighting the effectiveness of our relevant-paper-aware training approach. While GPT-4o tends to generate more positive comments, ReviewAgents, through stepwise analysis of strengths and weaknesses, produces rejection comments aligned with human reviewers’ assessments. We provide a complete case of the peer review comment generation process in the Appendix.
Figure 4: A Case Study. The green highlighted sections represent matching opinions, while the red highlighted sections represent opposing decisions.
6 Conclusion
In this paper, we aim to enhance the ability of LLMs to generate review comments, thereby advancing the automation of academic paper review. To achieve this objective, we collected papers and review comments and constructed the Review-CoT dataset, which includes the original papers, review comments with structured thinking processes, area chair comments, and relevant papers. Building on this, we employed a relevant-paper-aware method for fine-tuning LLMs and introduced ReviewAgents, a multi-agent review system that simulates the human review process, incorporating relevant paper retrieval, structured thinking, and a multi-role, multi-step approach. To comprehensively evaluate the capacity of LLMs to generate review comments, we developed ReviewBench, a benchmark comprising four testing dimensions. Experimental results demonstrate that our ReviewAgents surpass advanced LLMs, such as GPT-4o and Deepseek-R1.
7 Limitations
In this section, we discuss some of the limitations of our work.
7.1 Limited Research Area
Thanks to the openness of paper reviews in the AI field, our current work constructs datasets using review comments from AI conference papers. As a result, our approach to generating review comments is currently limited to the AI domain. In the future, we plan to collect papers and review comments from additional fields, thereby constructing datasets that encompass a broader range of disciplines, which will enhance the generalizability of our method across diverse research areas.
7.2 Disjoint Process for Review Generation
Our approach generates the final review comments through a multi-step generation process. During training, the reviewer agent and area chair agent are trained separately, which results in a disconnection between the training process and the inference phase. This may cause the input distribution for the area chair agent during the inference phase to differ from that during training. A potential improvement could involve performing joint training after separate supervised fine-tuning in both stages.
7.3 Limited Context Length
Due to the limitations imposed by the context window of large language models, ReviewAgents is constrained by a maximum paper length. Papers that exceed this length may experience truncation, leading to potential biases in the final review comments. Expanding the context window could result in increased training and inference costs. A potential solution to this issue is the batching of papers for processing.
Ethical Statement
This study constructs our dataset and benchmark using publicly available research paper data. All data used in this research is sourced from openly accessible academic repositories and databases, and no personal or sensitive information of authors or research participants is involved. Data processing and analysis adhered to the relevant terms of use of these data sources, and no privacy or confidentiality issues arose during the data collection process. The authors declare no conflicts of interest, and all analyses were conducted using automated methods, with complete transparency in the algorithms and data processing procedures. This research follows the ethical guidelines for the responsible use of data in academic research.
The potential risk of our research is the abuse of automatic peer review, which could lead to a decline in the quality of review comments and irresponsible feedback. We emphasize that our work is not intended to replace human reviewers. Instead, they serve as valuable tools for authors to self-edit their papers and as an aid for human reviewers. The results produced by the model should not be misconstrued as definitive and genuine comments on the respective papers.